OpenStudy (anonymous):

APEX GEOMETRY HELP! I need help finding the lenght of an arc, can someone help me?

OpenStudy (unklerhaukus):

the circumference of a circle is \(2\pi r\) the length of an arc over theta degrees is \[2\pi r \times \frac{\theta°}{360°}\]

OpenStudy (anonymous):

The way the circle os portrayed, its a triangle being made, the aec being the base, and the angle on top is 18 degrees

OpenStudy (unklerhaukus):

for example, if the angle of an arc is 90° \[L = 2\pi r \times \frac{90°}{360°}=2\pi r\times\frac14\]

OpenStudy (anonymous):

The opposite arc of arc im trying to find is 83 degrees

OpenStudy (unklerhaukus):

can you draw it?

OpenStudy (anonymous):

Thats all the information it gives me, it doesnt give me a radius or anything else

OpenStudy (anonymous):

Can we do that in mobile?

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):